OpenStudy (anonymous):

Convert 34/8 to a mixed number in simplest form

OpenStudy (anonymous):

How would u try to do this?

pooja195 (pooja195):

4 2/8

pooja195 (pooja195):

you divide 2 of the numbers denominator and numerator

hero (hero):

\[\frac{34}{8} = \frac{32 + 2}{8} = \frac{32}{8} + \frac{2}{8} = 4 \frac{2}{8} = 4 \frac{1}{4}\]

OpenStudy (anonymous):

4 1/4 actually....2 goes into 8 @pooja195

pooja195 (pooja195):

well yeah if it's in simplest form then it is 4 1/4

OpenStudy (anonymous):

yeah